If there are 7 periods in a school day, and a typical student takes 5 classes, how many different time patterns are possible for the student?
Answer by checkley71(8403)

the first subject can be in 1 of 7 periods, the seond can be in 1 of 6 periods etc.
5*7+4*6+3*5+2*4*1*3=35+24+15+8+3=85 different time patterns are possible.